# FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema

### v1

@@ 5 Conclusion
We present the Free-form Instruction-oriented Prompt Optimization (FIPO). The modular FIPO schema proposes to address Automatic Prompt Optimization as end-to-end text generation, flexibly taking raw task instruction, raw response and ground truth as inputs, for obtaining a new optimized instruction-oriented prompt. We hereby collect a large-scale, high-quality prompt preference dataset, employ with multiple fine-tuning strategies, and then validate the efficacy across five public benchmarks with various downstream generators.
@@
On the one hand, due to the current open-source models primarily concentrating on the “small-scale” around 10 billion parameters and otherwise “large-scale” around 70 billion parameters, we lack the opportunity to experiment with medium-sized models (e.g., 40 billion parameters). Based on the conclusions we have drawn, the 10B “small” LLMs are completely incapable of addressing free-from Automatic Prompt Optimization. Therefore, we would like to explore the critical model size threshold for this capability in our future research.
On the other hand, our current research on FIPO does not include optimization of the in-context examples, but rather focuses on prompt optimization for the task instructions. We maintain an open attitude towards related research.
Appendix A HyperParameters and Training Cost
We report hyperparameters and training cost in Table 6. We fine-tune all models on same computational node server equipped with 900G CPU RAM and 8 A100 40G GPUs. We only use single node for all experiments with 13B models, and LORA training of 70B models. As for full training of 70B models, we takes 4 nodes. While the batch size varies, we maintain a consistent global batch size on every node. In terms of different fine-tuning strategies, most hyperparameters are shared. The only difference lies on the learning rate, as we find larger learning rate would destroy the training stability of preference optimization. What’s more, all preference learning approaches relay on one external hyperparameter $\bm{\beta}$. We set its value as 0.01 empirically Tunstall et al. (2023).
@@ Appendix A HyperParameters and Training Cost
13B-Lora 13B-Full 70B-Lora 70B-Full
Nodes 1 1 1 4
Batch 16 8 2 2
Accumulations 1 2 8 8
@@
Figure 5: An overview of our dataset diversification step. It is recommended to view the details with colors.
Appendix B Dataset Diversification
We present our complete data diversification plans in Figure 5. We first collect response and prompt preference data as shown in the top. Afterwards, to better minimize the exposure gaps, we carry on the diversification combinations as shown within the table in the left middle of Figure 5. The diversification optionally selects raw instruction response $\bm{\hat{y}^{r}}$, ground truth response $\bm{y}^{r}$ and organization format. We thereby have eight various data format, and we demonstrate with two types in the bottom left and middle parts. The left corner one is an example format of type 3, in which only mandatory raw task instruction $\bm{x}^{r}$ is included. It worth to mention that this format type is also used during the testing phase as no responses are provided at that stage. Another type 6 located in the middle bottom transforms the initial generation format as a multi-choice one, and add a new binary raw instruction response $\bm{\hat{y}^{r}}$. Nevertheless, either type 3 or type 6 has fixed task description attached at the end.
The right side of Figure 5 illustrates specific incremented data for IPL fine-tuning. Equations 4 and  1 denote that only optimizer model $\bm{M_{o}}$ is included in preference optimization and supervised fine-tuning approaches. However, discrimination and instruction answering capacities are required in our iterative self-rewarding training, as shown in equations 7 and  8. Hence, we simply re-use the collected data in the top of Figure 5. 15k raw task instruction $\bm{x^{r}}$ and ground truth response $\bm{y^{r}}$ are paired as additional insturction following data of IPL, shown in the right middle part. In the meantime, we set a new meta prompt for the inner discrimination of IPL, shown in the right corner of Figure 5. Another 15k raw task instruction $\bm{x^{r}}$, ground truth response $\bm{y^{r}}$ and corresponding prompt labels $\bm{x^{o-}}$ and $\bm{x^{o+}}$ are re-used for data construction.
1:  Input require: Total number of iterations $\bm{N}$,
2:  Optimizer model $\bm{M_{o}}$,
3:  Discriminator model $\bm{M_{j}}$,
4:  Generator model $\bm{M_{g}}$,
5:  $\bm{M_{o}}\equiv\bm{M_{j}}\equiv\bm{M_{g}}$.
6:  Initilization: Raw Task Instruction $\bm{x^{r}}$,
7:  Raw Task Response $\bm{\hat{y}^{r}}$,
8:  Ground Truth Response $\bm{y}^{r}$,
9:  Suboptimal Optimized Prompt $\bm{x^{o-}}$,
10:  Optimal Optimized Prompt $\bm{x^{o+}}$.
11:  for $\bm{n}$ in $\bm{N}$ do
12:     if $\bm{n}$ > 1 then
13:        Get a prompt $\bm{x^{r+}}=\bm{M_{o}}(\bm{x^{r}},\bm{\hat{y}^{r}},\bm{y}^{r})$
14:        if $\bm{M_{j}}(\bm{x^{r}},\bm{y^{r}})\succeq\bm{M_{j}}(\bm{x^{r+}},\bm{y^{r}})$ then
15:           Reassign $\bm{x^{r+}}=\bm{x^{r}}$
16:        end if
17:        Get new response $\bm{\hat{y}^{r+}}=\bm{M_{g}}(\bm{x^{r+}})$
18:        Update raw task instruction $\bm{x^{r}}=\bm{x^{r+}}$
19:        Update raw task response $\bm{y}^{r}=\bm{\hat{y}^{r+}}$
20:     end if
21:     Update $\bm{M_{o}}$ with DPO or IPO loss.
22:  end for
Algorithm 1 Self-rewarding IPL Algorithm.
Appendix C Additional details of IPL
We first report the algorithmic narrative of IPL in Algorithm 1. In consistent with Equations 5,  6,  7 and  8, we first generate a new prompt $\bm{x^{r+}}$ with the optimizer model $\bm{M_{o}}$, then ask the model itself to judge if the newly optimized prompt is superior than the raw prompt $\bm{x^{r}}$. We takes the better prompt as the final one, and then generate a new task response $\bm{\hat{y}^{r+}}$ for next iteration. It should be noted that we start such self-rewarding updates after one epoch of warm-up.
We separately report representative accuracy of fine-tuned discrimination ability and proportions of more rigours samples involved in our IPL training pipeline in Table 7. We takes the fulling training of 70B model with IPO loss as the example, noted as “IPL-IPO-70B-eN (F)”, in which “-eN” refers to the “N”-th round of iteration. We can notice that the optimizer can easily handle the binary discrimination task with simultaneous training on the additional discrimination data. Nevertheless, the optimizer model itself gradually adds new superior prompt samples for dynamic optimization with a conservative attitude, as shown in the “Selection” column. The first epoch is for warm-up as aforementioned, therefore there are not scores reported.
Weighted Avg. Dis. Accuracy Selection
Human 47.80 N/A N/A
IPL-IPO-70B-e1 (F) 48.30 N/A N/A
IPL-IPO-70B-e2 (F) 50.81 100% 1.25%
IPL-IPO-70B-e3 (F) 52.13 100% 2.40%
Table 7: Analysis regards to details of IPL. We takes full training of 70B model with IPO loss as the example. The “-eN” notation refers to the “N”-th round of iteration in IPL training. The selection refers to the data proportion that be updated with newly generated prompts.
Appendix D Entire fine-tuning results
@@
Human 24.77 36.21 73.35 51.17 51.22 47.79
13B (L) 18.62 29.98 71.61 46.95 46.94 43.14
SFT 13B (F) 19.91 31.04 71.61 47.68 49.57 44.93
70B (L) 26.44($\uparrow$) 32.46 73.90($\uparrow$) 50.64 50.03 46.33
70B (F) 21.43 32.92 74.39($\uparrow$) 49.97 51.55($\uparrow$) 46.96
13B (L) 26.98($\uparrow$) 34.62 71.06 47.48 50.53 46.59
DPO 13B (F) 17.17 30.90 68.50 44.87 46.64 42.69
70B (L) 24.24 41.88($\uparrow$) 76.24($\uparrow$) 55.87($\uparrow$) 54.09($\uparrow$) 51.39($\uparrow$)
70B (F) 27.74(normal-$\uparrow$) 35.56 74.17($\uparrow$) 54.93($\uparrow$) 52.73($\uparrow$) 49.07($\uparrow$)
13B-e1 (L) 27.05($\uparrow$) 35.75 70.95 48.99 50.77 47.16
13B-e2 (L) 24.92($\uparrow$) 35.88 71.28 47.65 51.17 47.17
13B-e3 (L) 25.61($\uparrow$) 35.23 70.41 49.56 51.10 47.16
70B-e1 (L) 22.04 42.15($\uparrow$) 75.26($\uparrow$) 55.64($\uparrow$) 55.02($\uparrow$) 51.74($\uparrow$)
IPL-DPO 70B-e2 (L) 23.48 42.63($\uparrow$) 75.97($\uparrow$) 56.11($\uparrow$) 54.09($\uparrow$) 51.54($\uparrow$)
70B-e3 (L) 21.96 42.04($\uparrow$) 75.64($\uparrow$) 56.01($\uparrow$) 55.13($\uparrow$) 51.84($\uparrow$)
70B-e1 (F) 25.38($\uparrow$) 33.86 74.39($\uparrow$) 52.42($\uparrow$) 52.44($\uparrow$) 48.12($\uparrow$)
70B-e2 (F) 24.28 36.10 73.84($\uparrow$) 51.22($\uparrow$) 52.04($\uparrow$) 48.23($\uparrow$)
70B-e3 (F) 25.13($\uparrow$) 35.25 74.95($\uparrow$) 50.46 52.12($\uparrow$) 48.10($\uparrow$)
13B (L) 24.62 34.40 71.72 48.99 50.65 46.70
IPO 70B (L) 22.95 42.50($\uparrow$) 76.95($\uparrow$) 55.97($\uparrow$) 54.17($\uparrow$) 51.58($\uparrow$)
70B (F) 25.00($\uparrow$) 39.21($\uparrow$) 76.84($\uparrow$) 56.01($\uparrow$) 54.29($\uparrow$) 50.94($\uparrow$)
13B-e1 (L) 24.47 35.86 72.15 47.95 50.68 46.98
13B-e2 (L) 26.98($\uparrow$) 35.30 71.93 48.86 50.36 46.88
13B-e3 (L) 25.38($\uparrow$) 35.93 71.23 49.03 51.18 47.36
70B-e1 (L) 22.19 41.88($\uparrow$) 75.53($\uparrow$) 55.81($\uparrow$) 54.12($\uparrow$) 51.25($\uparrow$)
IPL-IPO 70B-e2 (L) 22.11 42.77($\uparrow$) 75.59($\uparrow$) 55.47($\uparrow$) 54.56($\uparrow$) 51.66($\uparrow$)
70B-e3 (L) 21.96 42.92(normal-$\uparrow$) 75.37($\uparrow$) 55.64($\uparrow$) 55.03($\uparrow$) 51.94($\uparrow$)
70B-e1 (F) 25.99($\uparrow$) 34.09 74.66($\uparrow$) 53.62($\uparrow$) 52.33($\uparrow$) 48.30($\uparrow$)
70B-e2 (F) 26.70($\uparrow$) 38.07($\uparrow$) 76.28($\uparrow$) 54.38($\uparrow$) 54.81($\uparrow$) 50.81($\uparrow$)
70B-e3 (F) 26.67($\uparrow$) 39.60($\uparrow$) 77.11(normal-$\uparrow$) 56.71(normal-$\uparrow$) 56.02(normal-$\uparrow$) 52.13(normal-$\uparrow$)
Table 8: Complete Evaluation results of downstream testing on out-of-box Tulu2-7B model. “Human” refers to use raw few-shot prompts written by humans, while other lines are tests with optimized prompts. The number attached with the benchmark name refers to the number of in-context examples (e.g., BBH (3) stands for 3-shot evaluation on BBH). The “(L)” or “(F)” letter attached with FIPO optimizer refer to parameter-efficient tuning LORA Hu et al. (2022) and full-parameter tuning, respectively. The “X” part of “IPL-X” refers to the different loss style mentioned in Equation 5. And the “-eN” notation refers to the “N”-th round of iteration in IPL training (e.g., 13B-e2 refers to the second iteration of fine-tuning the 13B model).
@@
We present entire fine-tuning results of our various experiments in Table 8. Apart from conclusions reported in Section 4.2, we further summarize several findings: (1) When fine-tuned with SFT approaches, 70B models perform better than 13B models as expected, so as to the full training compared with Lora. However, all SFT models perform worse than the human baseline; (2) In terms of two basic preference optimization training (DPO and IPO), we observe the Lora tuning performs better than the full-parameter training on the contrary. Nevertheless, 70B models are still superior to 13B models, and only the former ones obtains super-human results; (3) When it comes to our IPL approaches, there are differences when combined with DPO or IPO accordingly. IPL-DPO keeps a similar conclusion with DPO, as we find Lora tuning with 70B models behave the best, although our IPL strategy can improve with margins. As for IPL-IPO, we particularly find its full training works the best, and obtain an obvious growing trend. We suppose the regularized design in Equation 3 contributes to stable full-parameter improvements.
Appendix E Employed Meta Prompts
@@
We report all employed meta prompts in this section as reference for potential future researches. Please refer to Table 9,  10 and  11 for data collection meta prompt, data diversification meta prompt, and discrimination meta prompt, respectively.
@@
Table 9: The meta prompt used to harness data from GPT-3.5 and GPT-4 APIs. During harnessing, we replace placeholders “SP”, “SR” and “GR” with actual Raw Task Instruction $\bm{x^{r}}$, Raw Task Response $\bm{\hat{y}^{r}}$ and Ground Truth Response $\bm{y}^{r}$, respectively.
@@
Flexible meta prompt for optional raw task response
@@
Table 10: The meta prompt used for data diversification. During diversification, we replace placeholders “SP”, “SR” and “GR” with actual Raw Task Instruction $\bm{x^{r}}$, Raw Task Response $\bm{\hat{y}^{r}}$ and Ground Truth Response $\bm{y}^{r}$, respectively. Meanwhile, flexible prompts of optional raw task response and ground truth response are inserted by probability as introduced in Figure 3 and  5. Moreover, we add length suggestion with “GN” signal, using the length of optimal optimized prompt label $\bm{x^{o+}}$.
@@
Table 11: The meta prompt used for creating discrimination data and application of discrimination during IPL training. The creation of discrimination data has been introduced in the right part of Figure 5. In terms of inner discrimination of IPL, we replace placeholders “RP”, “PA”, “PB” and “GR” with actual Raw Task Instruction $\bm{x^{r}}$, Raw Task Instruction $\bm{x^{r}}$, Newly Optimized Prompt $\bm{x^{r+}}$ and Ground Truth Response $\bm{y}^{r}$, respectively.

### v2

@@ 5 Conclusion
We introduce FIPO, the Free-form Instruction-oriented Prompt Optimization. The modular FIPO template proposes to address APO as end-to-end text generation, flexibly taking naive prompt, naive response and ground truth as inputs, for obtaining a new optimized prompt. We hereby collect a large-scale prompt optimization preference dataset, employ with multiple fine-tuning strategies, and then validate the efficacy across objective benchmarks with various downstream generators.
@@
While FIPO demonstrates significant potential in optimizing prompts for various downstream tasks, there are several limitations to consider:
  • Overwhelmed Cheating Notes. As observed in the case analysis, FIPO sometimes provides overly detailed instructions that can be considered as “cheating notes”. This issue is particularly prevalent in tasks involving mathematical calculations. While this enhances performance, it may not align with the intended use of prompt optimization.
  • Evaluation Metrics. The current evaluation primarily focuses on accuracy metrics. While accuracy is important, other aspects such as the interpretability, fairness, and ethical implications of the optimized prompts should also be considered in future work.
  • Optimization of In-context Examplars. Our current research on FIPO does not include optimization of the in-context examples, but rather focuses on prompt optimization for the task instructions.
Appendix A Dataset Diversification
We present our complete data diversification plans in Figure 4. We first collect response and prompt preference data as shown in the top, according to pipeline introduced in section 2.4. Afterwards, to better minimize the exposure gaps, we carry on the diversification combinations as shown within the table in the left middle of Figure 4. The diversification optionally selects naive response $\bm{\hat{y}^{n}}$, ground truth response $\bm{y}^{n}$ and “generation” or “multi-choice” format. We thereby have eight various data format ($2\times 2\times 2$), and we demonstrate with two types in the bottom left and middle parts. The left corner one is an example format of type 3, in which only mandatory naive prompt $\bm{x}^{n}$ is included. It worth to mention that this format type is also used during the testing phase as no responses are provided at that stage. Another type 6 located in the middle bottom transforms the initial generation format as a multi-choice one, and add a new binary naive response $\bm{\hat{y}^{n}}$. Nevertheless, either type 3 or type 6 has same task description attached.
The right side of Figure 4 illustrates specific incremented data for IPL fine-tuning. Equations 11 and  8 denote that only optimizer model $\bm{M_{o}}$ is included in preference optimization and supervised fine-tuning approaches. However, discrimination and instruction answering capacities are required in our iterative self-rewarding training, as shown in equations 14. Hence, we simply re-use the collected data in the top of Figure 4. 15k naive prompts $\bm{x^{n}}$ and ground truth responses $\bm{y^{n}}$ are paired as additional instruction following data of IPL, shown in the right middle part. Meanwhile, we set a new meta prompt for the inner discrimination of IPL, shown in the right corner of Figure 4. Another 15k naive prompts $\bm{x^{n}}$, ground truth responses $\bm{y^{n}}$ and corresponding optimized prompts $\bm{x^{o-}}$ and $\bm{x^{o+}}$ are re-used for data construction.
Appendix B HyperParameters and Training Cost
We report hyperparameters and training cost in Table 7. We fine-tune all models on same computational node server equipped with 900G CPU RAM and 8 X A100 40G GPUs. We only use single node for fine-tuning 13B models, but use 4 nodes for 70B models. While the batch size varies, we maintain a consistent global batch size on every node. In terms of different fine-tuning strategies, most hyperparameters are shared. The only difference lies on the learning rate, as we find large learning rate for DPO and IPO will cause collapse. What’s more, all preference learning approaches relay on one external hyperparameter $\bm{\beta}$. We set its value as 0.01 empirically Tunstall et al. (2023).
@@ Appendix B HyperParameters and Training Cost
Tulu2 13B Tulu2 70B
Nodes 1 4
Batch 8 2
Accumulations 2 8
@@
Appendix C Entire fine-tuning results
@@
Naive 24.77 36.21 73.35 51.17 51.22 47.79
SFT 13B 19.91 31.04 71.61 47.68 49.57 44.93
70B 21.43 32.92 74.39 49.97 51.55 46.96
DPO 13B 17.17 30.90 68.50 44.87 46.64 42.69
70B 27.74 35.56 74.17 54.93 52.73 49.07
70B-e1 25.38 33.86 74.39 52.42 52.44 48.12
IPL-DPO 70B-e2 24.28 36.10 73.84 51.22 52.04 48.23
70B-e3 25.13 35.25 74.95 50.46 52.12 48.10
IPO 13B 18.42 33.55 72.03 49.14 48.20 44.93
70B 25.00 39.21 76.84 56.01 54.29 50.94
70B-e1 25.99 34.09 74.66 53.62 52.33 48.30
IPL-IPO 70B-e2 26.70 38.07 76.28 54.38 54.81 50.81
70B-e3 26.67 39.60 77.11 56.71 56.02 52.13
Table 8: Complete evaluation results of different fine-tuning strategies, tested by the same Tulu2-7B model. The “-eN” notation refers to the “N”-th round of iteration in IPL training (e.g., 13B-e2 refers to the second iteration of fine-tuning the 13B model).
@@
We present entire fine-tuning results of our various experiments in Table 8. Apart from conclusions reported in Section 3.2.2, we further summarize several findings: (1) When fine-tuned with SFT approaches, 70B models perform better than 13B models as expected. However, all SFT models perform worse than the naive baseline; (2) In terms of two basic preference optimization training (DPO and IPO), we observe that 70B models are still superior to 13B models, and only the former ones obtains super-human results; (3) When it comes to our IPL approaches, there are differences when combined with DPO or IPO accordingly. IPL-DPO keeps a similar conclusion with DPO. As for IPL-IPO, we particularly find its full training works the best, and obtain an obvious growing trend. We suppose the regularized design in Equation 10 contributes to stable improvements.
1:  Input require: Total number of iterations $\bm{E}$,
2:  Optimizer model $\bm{M_{o}}$,
3:  Initilization: Naive Prompt $\bm{x^{n}}$,
4:  Naive Response $\bm{\hat{y}^{n}}$,
5:  Ground Truth Response $\bm{y}^{n}$,
6:  for $\bm{e}$ in $\bm{E}$ do
7:     if $\bm{e}$ > 1 then
8:        New prompt $\bm{x^{n+}}=\bm{M_{o}}(\bm{x^{n}},\bm{\hat{y}^{n}},\bm{y}^{n})$
9:        if $\bm{M_{o}}(\bm{x^{n+}},\bm{y^{n}})\succ\bm{M_{o}}(\bm{x^{n}},\bm{y^{n}})$ then
10:           New response $\bm{\hat{y}^{n+}}=\bm{M_{o}}(\bm{x^{n+}})$
11:           Update $\bm{x^{n}}=\bm{x^{n+}}$
12:           Update $\bm{\hat{y}^{n}}=\bm{\hat{y}^{n+}}$
13:        end if
14:     end if
15:     Update $\bm{M_{o}}$ with DPO or IPO loss.
16:  end for
Algorithm 1 Self-rewarding IPL Algorithm.
Appendix D Additional details of IPL
We first report the algorithmic narrative of IPL in Algorithm 1. In consistent with Equations 12,  13, and  14, we first generate a new prompt $\bm{x^{n+}}$ with the optimizer model $\bm{M_{o}}$, then ask the optimizer also to judge if the newly optimized prompt is superior than the naive prompt $\bm{x^{n}}$. We takes the better prompt as the final one, and then generate a new task response $\bm{\hat{y}^{n+}}$ using the optimizer itself for next iteration. It should be noted that we start such self-rewarding updates after one epoch of warm-up.
As listed in the bottom of Table 8, IPL brings further benefits for IPO training. Compared with basic IPO preference tuning, to pump them into our IPL pipeline can advance to better accuracy, exhibiting a steeper upward growth of 2% per epoch.
we also report the accuracy of fine-tuned discrimination ability and proportions of more rigours samples involved in IPL in Table 9. We can notice that the optimizer can easily handle the binary discrimination task with simultaneous training on the additional discrimination data. We use 5% training data as the validation set, and find 100% classification accuracy. Based on this, the optimizer accurately update naive prompts $\bm{x}^{n}$ with self-generate new prompts $\bm{x}^{n+}$, introduced from line 9 to line 11 in Algorithm 1. Nevertheless, the optimizer model gradually update new superior prompt samples for dynamic optimization with a conservative attitude, as shown in the “Selection” column. We observe only 2.4% naive prompts and responses samples are upgraded in the end, while the self-game approach significantly promotes the optimization of downstream test prompts, as mentioned in Table 8.
Weighted Avg. Dis. Accuracy Selection
Naive 47.80 N/A N/A
IPL-IPO-70B-e1 48.30 N/A N/A
IPL-IPO-70B-e2 50.81 100% 1.25%
IPL-IPO-70B-e3 52.13 100% 2.40%
Table 9: Analysis regards to details of IPL. The “-eN” notation refers to the “N”-th round of iteration in IPL training. The selection refers to the data proportion that be updated with newly generated prompts. The first epoch is for warm-up, therefore there are not scores.
Figure 5: Analysis of overwhelmed cheating notes on 256 random optimized prompts for each benchmark.
Appendix E Overwhelmed Cheating Notes
We provide a proportion visualization of analyzing the possible over-optimized prompts in each testing benchmark, manually checking with 256 random samples, in Figure 5. The results show that mathematical questions in GSM8K and BBH are more likely to be added cheating notes of final answers, while other types of testing data receive moderate prompt optimization. Nevertheless, even taking into account this factor, the conclusion that our local optimizer improves the overall performance of general downstream generators by optimizing prompts reasonably does not change (§ 3.2.1).
Appendix F Employed Meta Prompts
@@
We report all employed meta prompts in this section as reference for potential future researches. Please refer to Table 10,  11 and  12 for data collection meta prompt, data diversification meta prompt, and discrimination meta prompt, respectively.
@@
Table 10: The meta prompt used to harness data from GPT-3.5-turbo and GPT-4 APIs. During harnessing, we replace placeholders “SP”, “SR” and “GR” with actual naive prompt $\bm{x^{n}}$, naive response $\bm{\hat{y}^{n}}$ and ground truth response $\bm{y}^{n}$ from each seed data sample, respectively.
@@
Flexible meta prompt for optional naive response
@@
Table 11: The meta prompt used for data diversification. During diversification, we replace placeholders “SP”, “SR” and “GR” with actual naive prompt $\bm{x^{n}}$, naive response $\bm{\hat{y}^{n}}$ and ground truth response $\bm{y}^{n}$, respectively. Meanwhile, flexible prompts of optional naive response and ground truth response are inserted by probability as introduced in Figure 3 and  4. Moreover, we add length suggestion with “GN” signal, using the length of chosenl optimized prompt label $\bm{x^{o+}}$.
@@
Table 12: The meta prompt used for creating discrimination data and application of discrimination during IPL training. The creation of discrimination data has been introduced in the right part of Figure 4. In terms of inner discrimination of IPL, we replace placeholders “RP”, “PA”, “PB” and “GR” with naive prompt $\bm{x^{n}}$, naive prompt $\bm{x^{n}}$, newly optimized prompt $\bm{x^{n+}}$ and ground truth response $\bm{y}^{n}$, respectively.